# Mamba State-Space Models Are Lyapunov-Stable Learners

### v2

@@ Appendix D Experimental Details
Appendix E Mamba-2 mixed-precision inference pretrained LLM performance tables
Table 3: Mean full-precision (FP32) divergence in MMLU performance for mixed-precision inference. Divergence is averaged over {0, 1, 3, 5}-shot performance. Pretrained checkpoints are used for Mamba (M), Mamba-2 (M2), Pythia (P), OLMo [19] (O), and Phi-1.5 [33] (Phi) models.
Model M M2 P M M2 P M M2 P O M M2 P Phi M M2 P
Size 130m 130m 160m 370m 370m 410m   790m 780m 1b 1.4b 1.3b 1.4b 1.5b 2.8b 2.7b 2.8b
FP16 $\mu$ 0.03 0.07 0.35 0.05 0.05 0.06 0.21 0.12 0.05 0.04 0.04 0.04 0.07 0.03 0.15 0.26 0.12
BF16 $\mu$ 0.05 0.67 1.45 0.20 0.52 0.20 0.66 0.29 0.16 0.13 0.31 0.40 0.13 1.05 1.17 1.02 0.11
Appendix F Adapted PEFT-layer effects on Mamba-2 zero-shot performance
As described in Section 3.2, both Mamba and Mamba-2 store the majority of time-varying SSM parameters in a large memory buffer (instantiated as a PyTorch linear layer’s weight matrix), denoted as $\mathbf{W}$. For PEFT via LoRA, we ablate the impact of targeting $\mathbf{W}$, as well as the impact of Theorem 3, where time-varying variables accessed through $\mathbf{W}$ were shown to be regularized through weight tying.
Firstly, we instruction tune Mamba-2 models using LoRA by both adapting $\mathbf{W}$ and adapting all linear layers other than $\mathbf{W}$. Mamba-2 MambaBlocks contain two linear layers, thus the former targets only the $\mathbf{W}$ linear layer whereas the latter targets the other linear layer in each block. As shown in Table 4, PEFT with only $\mathbf{W}$ targeted within the MambaBlock near uniformly results in better performance than only targeting other layers, with the former outperforming the latter on 32 out of 35 natural language tasks. This makes intuitive sense since $\mathbf{W}$ represents the majority of Mamba’s time-varying parameters (as previously mentioned). Thus, while previous works [21, 11] have displayed the importance of Mamba’s time-varying parameters for pretraining performance, this verifies the importance of Mamba’s time-varying parameters for instruction tuning performance.
Table 4: Zero-shot performance for instruction tuned Mamba-2 models where: denotes the large memory buffer $\mathbf{W}$ containing the majority of temporal variables (described in Section 3.2) is targeted for LoRA adaptation, and denotes $\mathbf{W}$ is not adapted. The top-performance for each task per model is highlighted in bold .
Model $\mathbf{W}$ targeted? LAMBADA LAMBADA HellaSwag PIQA Arc-E Arc-C WinoGrande
ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$
Mamba-2 130M 15.37 45.16 35.41 65.02 47.94 24.91 52.17
16.95 43.30 35.24 64.85 47.47 24.06 52.17
Mamba-2 370M 8.03 54.16 46.87 69.53 53.49 27.73 57.14
8.50 53.74 46.59 70.67 54.80 26.79 55.64
Mamba-2 780M 5.79 61.61 55.14 71.98 61.11 29.18 60.09
5.86 61.63 54.94 72.03 60.98 28.41 60.22
Mamba-2 1.3B 4.54 65.73 60.88 73.67 66.16 34.56 61.80
5.05 65.44 59.84 73.39 64.10 33.11 61.09
Mamba-2 2.7B 4.05 69.63 66.73 76.50 69.95 36.77 64.56
4.10 69.61 66.60 76.39 69.53 36.26 63.93
Next, we present a counterintuitive experiment to demonstrate the impact of Theorem 3. As before, we instruction tune Mamba-2 models using LoRA, focusing on the two largest model sizes. This time, however, we target all linear layers within Mamba-2’s MambaBlock and contrast this with targeting only $\mathbf{W}$ within the MambaBlock. Displayed in Table 5, we can see that adapting more SSM parameters does not necessarily lead to improved performance across the board. Rather, adapting only $\mathbf{W}$ for Mamba-2 outperforms adapting all linear layers on eight of the 14 natural language tasks.
The results in Table 5 thus present two cases: (a) regularization: only adapting parameters which lead to weight tying (by Theorem 3) and (b) increased learning capacity: adapting more parameters, at the cost of learning unregularized parameters. Through regularization via weight tying [42], the former leads to good generalization. In contrast, while adapting more parameters, (b) includes unregularized variables, leading to positive improvements on only a minority of tasks when compared to the fine-tuning of fewer, but regularized, parameters. We note that adapting all linear layers is often performed when fine-tuning attention-based LLMs using LoRA [12]. Thus, a similar regularization-vs-capacity tradeoff for Transformers LLMs may present itself by carefully studying LoRA’s affects when targeting specific linear layers in attention-based architectures.
Table 5: Zero-shot performance for instruction tuned Mamba-2 models. denotes only $\mathbf{W}$ was targeted for LoRA adaptation within the MambaBlock, and thus denotes both linear layers within each MambaBlock were adapted. The top-performance for each task per model is highlighted in bold.
Model Only $\mathbf{W}$ targeted? LAMBADA LAMBADA HellaSwag PIQA Arc-E Arc-C WinoGrande
ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$
Mamba-2 1.3B 4.54 65.73 60.88 73.67 66.16 34.56 61.80
4.94 62.95 60.06 73.78 64.02 34.81 61.33
Mamba-2 2.7B 4.05 69.63 66.73 76.50 69.95 36.77 64.56
3.77 68.39 67.51 75.95 69.53 38.31 65.35

### v3

@@ Appendix F Experimental Details
All model checkpoints were evaluated on all benchmarks and few-shot settings using the LM evaluation harness from Eleuther AI (Gao et al., 2023), version 0.4.2. Pythia and Mamba Huggingface checkpoints were used for all inference and fine-tuning experiments, e.g., EleutherAI/pythia-160m and state-spaces/mamba-130m-hf for the smallest respective models. All fine-tuning experiments were run using package versions Transformers 4.40.0.dev0, Accelerate 0.28.0, TRL 0.8.1, PyTorch 2.2.1+cu121, and PEFT 0.10.0. All Mamba-2 models were run using mamba-ssm v2.2.2 using Huggingface checkpoints, e.g., state-spaces/mamba-130m for the smallest model.
For MPFT, Flash Attention 2.0 (Dao et al., 2022) via flash_attn 2.5.7 was used for Pythia models. For FP16 and BF16 inference results, Flash Attention 2.0 was used for both Pythia and OLMo models. For OLMo results, the 336B-token checkpoint was used by specifying revision=step80000-tokens336B.
All Alpaca and OpenHermes fine-tuning experiments used the following training recipe (adapted from (Tunstall et al., 2023)): AdamW_torch optimizer, cosine annealing schedule, no gradient accumulation, maximum norm of 1.0 for gradient clipping, and no warmup steps. Training epochs used for all Alpaca and OpenHermes experiments were three and one, respectively. For both Pythia and Mamba models, the learning rate and LoRA dimension $r$ were scaled to improve performance of smaller models (per-model values listed in Table 1).
For SLL LoRA, targeted Mamba layers were {x_proj, embeddings, in_proj, out_proj}; x_proj is a large MambaBlock memory buffer which, when targeted by LoRA, adapts parameters $\bm{\Delta}_{t},\mathbf{B}_{t}$, and $\mathbf{C}_{t}$. Pythia targeted SLL LoRA layers were {dense, embed_in, query_key_value, dense_h_to_4h,dense_4h_to_h}, chosen to balance performance across model sizes.
@@
Table 2: Pretrained model performance. Model checkpoints were evaluated on all benchmarks and few-shot settings using the LM evaluation harness from Eleuther AI (Gao et al., 2023). LAMBADA zero-shot is more effective for the model sizes considered (further discussed in (Xie et al., 2021; Brown et al., 2020)) and thus excluded from few-shot performance averages. Highlighted in bold is the top-performing few-shot learner per benchmark and model grouping.
Model NN-shot LAMBADA LAMBADA HellaSwag PIQA Arc-E Arc-C WinoGrande 0-shot incr.
ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ Mean % $\uparrow$
Mamba 130M 0 16.0 44.3 35.2 64.7 48.0 24.3 52.6
1 19.3 38.2 35.1 64.6 47.0 23.5 50.8 -1.9
3 23.1 35.2 35.0 65.1 49.1 24.0 51.0 -0.4
5 24.4 36.2 34.9 64.9 49.1 23.7 50.0 -1.2
Mamba-2 130M 0 16.8 43.9 35.3 64.9 47.4 24.2 52.6
1 20.6 37.9 34.9 64.1 46.9 23.1 51.3 -2.0
3 24.3 35.1 34.9 64.4 49.0 24.7 52.9 0.9
5 26.5 34.9 34.6 64.4 48.6 24.8 51.7 0.2
Pythia 160M 0 38.2 32.7 30.2 61.8 43.4 23.8 51.0
1 47.2 28.2 30.6 62.2 43.4 23.7 49.3 -0.4
3 63.7 24.7 30.5 61.9 44.8 22.9 51.3 0.1
5 66.3 25.3 30.4 62.6 43.4 23.1 50.8 -0.2
Mamba 370M 0 8.1 55.6 46.5 69.5 54.9 27.8 55.3
1 9.7 49.8 45.9 69.3 57.4 26.5 54.7 -0.5
3 10.9 48.4 46.2 69.5 58.8 28.4 53.8 1.2
5 11.4 48.6 46.2 69.4 58.3 28.0 55.9 1.5
Mamba-2 370M 0 8.0 55.9 46.9 70.5 54.8 26.7 55.4
1 9.8 50.3 46.4 70.5 56.5 26.8 54.2 0.0
3 11.3 48.5 46.6 70.2 59.0 26.9 54.3 1.0
5 12.5 46.6 46.7 70.3 58.5 28.2 53.3 1.5
Pythia 410M 0 10.8 51.5 40.6 66.9 52.0 24.1 53.4
1 12.3 47.1 40.5 68.0 53.8 25.6 52.4 1.8
3 14.4 43.2 40.9 67.9 55.1 26.9 54.0 4.2
5 14.6 44.1 40.8 68.1 54.6 26.6 53.4 3.5
Mamba 790M 0 6.0 61.4 55.1 72.3 61.2 29.5 55.9
1 7.1 55.9 54.5 72.4 63.0 30.2 56.9 1.3
3 8.1 54.5 54.2 72.3 63.5 31.4 57.1 2.2
5 8.8 52.9 54.6 72.6 64.4 31.9 57.2 3.1
Mamba-2 780M 0 5.9 61.7 54.9 72.0 61.0 28.5 60.2
1 7.1 55.5 54.7 72.4 62.3 32.1 57.1 1.9
3 8.6 53.3 54.7 72.5 62.8 32.3 57.8 2.5
5 9.9 51.4 55.2 72.1 62.8 32.2 56.8 2.2
Pythia 1B 0 7.9 56.3 47.2 70.7 57.0 27.0 53.4
1 8.0 51.8 47.3 70.7 57.1 28.2 53.4 1.0
3 10.5 48.2 47.5 71.2 59.2 28.0 54.3 2.2
5 10.9 48.4 47.3 71.4 58.7 28.4 53.1 1.9
Mamba 1.4B 0 5.0 64.9 59.2 74.1 65.5 32.9 61.3
1 5.8 60.6 58.2 74.7 64.5 33.0 60.9 -0.5
3 6.6 58.9 58.9 73.6 66.1 34.5 60.9 0.7
5 7.0 58.3 59.0 74.1 66.4 35.5 60.4 1.5
Mamba-2 1.3B 0 5.0 65.6 60.0 73.2 64.2 33.1 61.1
1 6.0 60.1 59.4 73.1 65.6 35.3 59.4 1.0
3 6.7 58.6 60.1 73.4 66.5 35.4 61.9 2.5
5 7.0 58.6 60.2 73.7 66.5 35.9 61.4 2.7
Pythia 1.4B 0 6.1 61.7 52.1 70.9 60.5 28.5 57.4
1 7.0 56.3 52.1 71.4 62.0 29.5 57.5 1.4
3 7.9 54.4 52.6 70.9 63.9 31.1 56.8 2.9
5 8.0 54.4 52.8 71.0 63.2 31.3 57.8 3.3
Mamba 2.8B 0 4.2 69.1 66.1 75.2 69.6 36.4 63.3
1 5.0 63.7 65.6 75.6 69.9 37.1 63.9 0.6
3 5.5 62.8 65.5 75.3 70.8 38.1 65.1 1.7
5 5.7 62.5 66.1 76.1 70.9 38.1 64.6 2.0
Mamba-2 2.7B 0 4.1 69.6 66.6 76.4 69.5 36.3 63.9
1 4.8 65.1 65.9 75.1 70.0 38.6 65.1 1.3
3 5.3 63.9 66.8 75.2 71.9 41.0 64.1 3.1
5 5.7 62.3 67.1 75.3 70.7 41.2 65.9 3.6
Pythia 2.8B 0 5.0 64.7 59.3 73.9 64.2 32.9 59.8
1 5.7 60.9 59.4 73.8 66.8 34.8 59.0 1.7
3 6.2 59.1 59.9 74.7 67.4 34.9 60.8 2.9
5 6.5 59.1 60.2 74.5 67.1 35.0 61.3 3.1
Table 3: Instruction tuned model performance. Model checkpoints were evaluated on all benchmarks and few-shot settings using the LM evaluation harness from Eleuther AI (Gao et al., 2023). LAMBADA zero-shot is more effective for the model sizes considered (further discussed in (Xie et al., 2021; Brown et al., 2020)) and thus excluded from few-shot performance averages. Highlighted in bold is the top-performing few-shot learner per benchmark and model grouping.
Model NN-shot LAMBADA LAMBADA HellaSwag PIQA Arc-E Arc-C WinoGrande 0-shot incr.
ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ Mean % $\uparrow$
Mamba 130M 0 12.9 46.5 35.1 64.2 48.7 25.5 51.7
1 17.8 38.1 35.0 64.2 48.6 24.9 52.2 -0.4
3 22.3 35.3 34.8 64.2 50.2 24.5 50.6 -0.8
5 23.6 35.9 34.7 64.7 49.8 24.6 50.2 -0.9
Mamba-2 130M 0 15.2 44.5 35.1 64.5 47.2 24.7 52.2
1 21.9 36.1 34.5 64.3 46.8 24.0 50.8 -1.7
3 26.9 33.3 34.7 65.1 48.5 25.2 51.5 0.6
5 29.0 33.8 34.5 64.8 48.7 25.1 51.3 0.4
Pythia 160M 0 30.2 36.1 30.0 62.2 44.7 23.6 50.3
1 44.5 29.1 30.4 62.0 44.0 23.6 50.5 -0.0
3 66.7 25.5 30.3 62.8 45.2 22.8 49.8 -0.3
5 70.4 25.3 30.5 62.9 44.1 23.4 50.8 0.3
Mamba 370M 0 7.2 56.0 46.3 69.2 55.3 27.7 56.0
1 9.3 49.9 45.7 68.7 57.1 28.3 55.4 0.5
3 10.4 49.4 45.7 68.9 58.7 29.7 54.1 1.6
5 11.0 48.3 45.7 70.1 59.3 29.1 54.5 1.9
Mamba-2 370M 0 7.6 54.7 46.8 69.3 52.2 27.0 56.0
1 9.9 48.3 46.0 69.6 55.7 28.8 55.2 2.1
3 11.8 46.3 46.3 70.1 59.0 29.1 54.5 3.6
5 12.6 45.5 46.3 70.8 59.6 29.5 53.0 3.8
Pythia 410M 0 13.3 46.4 40.9 67.4 52.7 25.4 53.4
1 17.2 40.4 40.5 68.4 53.6 25.7 53.0 0.5
3 21.1 37.4 40.9 67.7 55.7 27.1 52.6 2.3
5 21.5 38.2 40.7 67.8 55.7 27.3 53.8 2.8
Mamba 790M 0 5.2 62.8 55.6 72.8 62.4 30.6 56.2
1 6.3 56.6 54.9 72.7 64.6 31.7 56.3 1.2
3 7.0 55.6 54.7 72.4 65.3 33.2 57.5 2.7
5 7.5 54.6 54.9 72.9 65.6 33.8 57.2 3.2
Mamba-2 780M 0 4.9 63.4 55.8 71.7 61.1 30.6 59.2
1 6.6 55.2 54.4 72.7 64.2 34.0 57.6 2.5
3 7.8 52.7 54.9 73.5 65.0 34.6 57.8 3.6
5 8.6 52.8 54.8 73.4 64.6 34.0 58.0 3.1
Pythia 1B 0 7.7 56.6 47.3 70.8 57.1 26.7 53.4
1 8.8 52.0 47.4 70.7 57.5 28.8 53.6 1.8
3 10.2 48.7 47.5 71.4 59.0 28.5 54.4 2.6
5 10.6 48.8 47.4 71.5 58.9 28.4 53.0 2.0
Mamba 1.4B 0 4.6 64.8 59.3 74.3 65.2 35.1 62.3
1 5.4 60.3 58.2 74.3 66.7 35.7 62.8 0.6
3 6.1 59.3 58.4 74.1 67.4 36.6 61.8 1.0
5 6.3 58.8 58.8 74.5 68.3 37.0 59.9 1.1
Mamba-2 1.3B 0 4.9 63.0 60.1 73.8 64.0 34.8 61.3
1 6.1 58.2 59.2 74.2 67.0 35.0 60.1 0.5
3 7.0 56.6 59.4 73.7 67.8 36.6 59.9 1.5
5 7.2 56.5 59.9 73.5 68.5 36.7 60.7 2.2
Pythia 1.4B 0 5.2 63.6 52.9 71.1 61.2 30.3 58.2
1 6.2 57.4 52.7 71.7 62.2 30.6 56.9 0.2
3 7.0 56.1 53.1 71.1 64.5 32.8 56.8 2.3
5 7.1 55.5 53.3 71.2 63.8 33.5 57.5 2.9
Mamba 2.8B 0 4.0 67.7 66.4 75.6 68.4 36.6 64.2
1 4.8 63.3 65.9 76.2 70.9 39.4 64.6 2.4
3 5.3 62.1 65.7 75.8 71.3 39.1 65.4 2.4
5 5.4 61.9 66.2 77.2 71.4 40.4 66.1 3.9
Mamba-2 2.7B 0 3.8 68.4 67.5 76.0 69.5 38.3 65.3
1 4.5 63.8 66.7 76.0 71.8 41.5 67.1 2.6
3 5.0 62.3 67.3 76.2 73.3 44.4 66.0 4.5
5 5.3 61.8 67.4 76.4 72.4 44.5 65.0 4.1
Pythia 2.8B 0 5.0 64.7 59.3 74.0 64.7 33.3 59.2
1 5.6 60.8 59.5 74.0 66.7 34.9 59.3 1.7
3 6.1 59.2 59.9 75.0 67.5 34.9 60.9 2.9
5 6.5 59.0 60.4 74.5 67.0 35.1 61.2 3.0
F.1 Hardware throughput and memory-utilization improvements given PEFT and MPFT
With stable dynamics and observed divergences smaller than comparable Transformers, we show that MPFT and PEFT may be used to significantly increase GPU-training throughput for Mamba SSMs. To demonstrate such improvements, we utilize the previous fine-tuning settings for the Alpaca dataset. However, we now adjust the batch size to maximize throughput per MPFT and PEFT configuration.
For each MPFT and PEFT configuration, the average tokens-per-second (ATPS) is calculated as the total tokens used for fine-tuning divided by total training time, and the maximum memory-per-token (MMPT) is calculated as the maximum GPU memory utilization incurred (over the entire fine-tuning run) divided by the total number of tokens in each mini-batch. Results are plotted in Figure 7.
Both throughput and memory utilization improve as the number of Mamba parameters increases in Figure 7. Compared to the full-precision full fine-tuning of Mamba 790M (the largest model supported by an A10G’s memory capacity), evaluated MPFT and PEFT combinations result in an average 2.15 times more training tokens-per-second while reducing per-token memory utilization by an average 62.7%. Across all model sizes, evaluated MPFT and PEFT combinations result in an average 1.74 times more training tokens-per-second while reducing per-token memory utilization by an average 47.2% compared to respective full-precision fine-tuned runs. Furthermore, while full fine-tuning is no longer possible on a single A10G for Mamba models greater than 790 million parameters, MPFT and PEFT allow training Mamba models up to 2.8 billion parameters on GPUs with as little as 24 GB onboard memory.
(a) Average tokens-per-second $\uparrow$.
(b) Maximum memory-per-token $\downarrow$.
Appendix G ICL as an emergent ability of Mamba SSMs
We study the emergent behavior (as a function of model size) of Mamba/Mamba-2 SSMs’ ICL abilities on natural language tasks by comparing to a larger number of Transformer-based LLMs of varying sizes. We compare to OpenELM (Mehta et al., 2024) (sizes 270M, 450M, and 1.1B), TinyLlama 1.1B (Zhang et al., 2024), and OLMo 1.2B (Groeneveld et al., 2024). To limit the emergent effects on both parameter size and pretraining token counts, we did not evaluate models greater than 2.8 billion parameters and chose open-source checkpoints as close as possible to the 300 billion total pretraining tokens used for Mamba, Mamba-2, and Pythia models. Thus, pretraining token counts for OpenELM, TinyLlama, and OLMo models were 429 billion, 503 billion, and 336 billion, respectively. We note this potentially biases ICL performance in favor of the newly evaluated Transformer-based LLMs, and that direct comparisons between Mamba, Mamba-2, and Pythia are the most fair (as these three classes of models were all pretrained on the same dataset for the same number of total pretraining tokens).
Instruction tuning greatly smooths ICL performance across both parameter classes. While instruction tuned SSMs and Transformers of 160 million parameters or fewer continue to display slight or detrimental ICL abilities, all parameters of 270 million and greater show positive ICL abilities. For instruction tuned models of greater than 450 million parameters, all SSMs and Transformers show positive ICL abilities, with Mamba-2 2.7B greatly outperforming all other models (both SSM and Transformer) in this class.
Thus, in terms of ICL as a function of SSM model size, while no clear trend presents itself for pretrained models, ICL emerges for instruction tuned Mamba and Mamba-2 SSMs of size 370 million and greater.
(a) $450\text{M}<\text{Parameter total}$, Pretrained
(b) $450\text{M}<\text{Parameter total}$, ALL LoRA
(c) $\text{Parameter total }\leq 450\text{M}$, Pretrained
(d) $\text{Parameter total }\leq 450\text{M}$, ALL LoRA
Figure 8: Instruction tuning improves Mamba-2 ICL performance past Transformer LLMs. ALL LoRA models were instruction fine-tuned on the OpenHermes dataset for one epoch. Performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot over five standard natural language benchmarks: HellaSwag, PIQA, Arc-E, Arc-C, and WinoGrande. Mamba* and Mamba’ refer to recent Mamba-1 PEFT methods, i.e., $\text{LoRA}_{p}(X)$ and Additional-Scan, respectively, from Yoshimura et al. (2025).
Appendix H Expanded divergence results: Alpaca and LIMA fine-tuning, MMLU and Winogrande benchmarks, Mean and Standard Deviation divergences
We extend the non-divergent Mamba fine-tuning results from Section 4. Recall that the following MPFT and PEFT configurations are considered to fine-tune each considered LLM:
  1. 1.
    Full fine-tuning in FP32
  2. 2.
    Full fine-tuning in FP16
  3. 3.
    Full fine-tuning in BF16
  4. 4.
    ALL LoRAfine-tuning in FP32
  5. 5.
    ALL LoRAfine-tuning in FP16
  6. 6.
    ALL LoRAfine-tuning in BF16
  7. 7.
    SLL LoRAfine-tuning in FP32
  8. 8.
    SLL LoRAfine-tuning in FP16
  9. 9.
    SLL LoRAfine-tuning in BF16
In addition to the Alpaca dataset (Taori et al., 2023), we also fine-tune all models using the LIMA dataset (Zhou et al., 2024). Models are trained using LIMA for 5 epochs, while all other settings follow the fine-tuning recipe used for Alpaca (described in Appendix F).
For natural language benchmarks, in addition to MMLU, we evaluate each fine-tuned model using Winogrande (Sakaguchi et al., 2021). Recall that, for each benchmark, divergence between a mixed-precision fine-tuned model is measured between its full-precision counterpart and averaged over {0, 1, 3, 5}-shot performance. In addition to the average divergence, we also include the standard deviation of divergence. Thus, in total, 144 LLMs were fine-tuned, 576 MMLU evaluations were conducted, and 576 Winogrande evaluations were conducted.
Figure 9 displays results for Alpaca and MMLU, Figure 10 displays results for Alpaca and Winogrande, Figure 11 displays results for LIMA and MMLU, Figure 12 displays results for LIMA and Winogrande. Summary statistics for all experiments are presented in Table 4. While OpenELM exhibits large deviation spikes for both Alpaca benchmark evaluations–and Pythia exhibits large deviation spikes for all four evaluations–Mamba does not exhibit a single large deviation spike on any benchmark for all considered model sizes and MPFT/PEFT configurations (i.e., 18 total configurations excluding the full-precision baselines). Furthermore, Mamba models are significantly more stable for MPFT/PEFT compared to Transformer-based LLMs. E.g., for MMLU evaluations, Alpaca fine-tuning with Mamba models is an average 2.6 times smaller in mean divergence than both Pythia and OpenELM models, while LIMA fine-tuning with Mamba models is an average 7 and 3.25 times smaller in mean divergence than Pythia and OpenELM models, respectively.
Table 4: Summary of divergence results for Alpaca and LIMA fine-tuning datasets, MMLU and Winogrande benchmarks, and Mamba, OpenELM, and Pythia models. For each deviation summary statistic per fine-tuning dataset and benchmark, the lowest deviation is highlighted in bold.
(Fine-tuning dataset), Benchmark Architecture Large deviation spikes $\downarrow$ Avg mean divergence $\downarrow$ Std mean divergence $\downarrow$
(Alpaca, MMLU ) Pythia 1 0.37 0.41
OpenELM 1 0.37 0.32
Mamba 0 0.14 0.08
(Alpaca, Winogrande ) Pythia 4 0.72 0.58
OpenELM 3 0.59 0.37
Mamba 0 0.25 0.09
(LIMA, MMLU ) Pythia 1 0.28 0.34
OpenELM 0 0.13 0.15
Mamba 0 0.04 0.03
(LIMA, Winogrande ) Pythia 3 0.45 0.45
OpenELM 0 0.36 0.18
Mamba 0 0.11 0.12
(a) Mean full-precision (FP32) divergence in MMLU performance.
Figure 9: Alpaca fine-tuning, MMLU evaluation. Mamba, Pythia, and OpenELM models are fine-tuned over the Alpaca dataset using different combinations of MPFT and PEFT. Full fine-tuning (i.e., no PEFT adapters) is denoted as Full.
(a) Mean full-precision (FP32) divergence in Winogrande performance.
Figure 10: Alpaca fine-tuning, Winogrande evaluation. Mamba, Pythia, and OpenELM models are fine-tuned over the Alpaca dataset using different combinations of MPFT and PEFT. Full fine-tuning (i.e., no PEFT adapters) is denoted as Full.
(a) Mean full-precision (FP32) divergence in MMLU performance.
Figure 11: LIMA fine-tuning, MMLU evaluation. Mamba, Pythia, and OpenELM models are fine-tuned over the LIMA dataset using different combinations of MPFT and PEFT. Full fine-tuning (i.e., no PEFT adapters) is denoted as Full.
(a) Mean full-precision (FP32) divergence in Winogrande performance.
Figure 12: LIMA fine-tuning, Winogrande evaluation. Mamba, Pythia, and OpenELM models are fine-tuned over the LIMA dataset using different combinations of MPFT and PEFT. Full fine-tuning (i.e., no PEFT adapters) is denoted as Full.